The area of a trapezoid is 18sq. ft and the sum of its bases is 6 ft. Find the area of a square whose side is equal to the height of the trapezoid.
Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

Area = (height) * (sum of bases)/2
18 = h * 3/2
h = 12

square area = 12^2 = 144 ft^2

Answer by ikleyn(53765)

.
The area of a trapezoid is 18sq. ft and the sum of its bases is 6 ft. Find the area of a square whose side is equal to the height of the trapezoid.
~~~~~~~~~~~~~~~~~~


            The solutions by   @CubeyThePenguin  is INCORRECT.

            I came to bring you the correct solution.


Area = (height) * (sum of bases)/2


    18 = h * 6/2


    h = 18/3 = 6


square area = 6 * 6 = 36 ft^2.     ANSWER

~~~~~~~~~~~~~~~~

Hello, @CubeyThePenguin, it looks like you are in bad condition today.

Consider to stop your job at the forum for some time and take a break.